# Zihan Tan    Suyuan Huang    Guancheng Wan    Wenke Huang    He Li    Mang Ye

### v2

@@ 5.2 Experiment Results
In this section, we comprehensively evaluate the proposed $S^{2}$ by addressing the following questions:
  • Q1: How does $S^{2}$ perform compared to existing methods in subgraph-FL?
  • Q2: What is the impact of each component of $S^{2}$ on its performance?
  • Q3: Does $S^{2}$ exhibit good stability?
  • Q4: Does the effectiveness of NLIR help alleviate the impact of SIS decline in subgraph-FL?
Q1:How does $S^{2}$ perform compared to existing methods in subgraph-FL?
We present the results of node classification tasks across various FGL scenarios using multiple graph datasets, and we summarize the final average test accuracy in Tab. 1. Our proposed method, $S^{2}$, demonstrates superior performance by outperforming all other baseline approaches in five out of the six datasets. This consistent superiority highlights the effectiveness of $S^{2}$ in handling diverse graph structures and data distributions inherent to different datasets.
Q2:What is the impact of each component of $S^{2}$ on its performance?
To evaluate the individual contributions of the proposed NLIR and FGMA strategies within the $S^{2}$ framework, we conducted ablation experiments on the Cora and Citeseer datasets. In this study, we systematically removed each component to assess its impact on model performance. The results of the ablation study are presented in Tab. 2 which demonstrate that both NLIR and FGMA independently contribute to the overall performance.
@@
Table 2: Ablation study of key components (NLIR, FGMA) of $S^{2}$ on Cora and Citeseer datasets.
Q3: Does $S^{2}$ exhibit good stability?
We evaluated the stability and adaptability of our approach on the Cora and Citeseer datasets through separate hyperparameter tuning and client partitioning experiments. Specifically, we varied the scaling factors of NLIR and FGMA to investigate how different parameter settings influence the model performance and robustness. Additionally, we conducted experiments with varying client scales
Varying Scaling Factors of NLIR and FGMA. For the NLIR method, we test scaling factors set to 100, 50, 10, and 1. For the FGMA method, the settings were 0.01, 0.05, 0.5, and 1. The results, shown in Fig. 4, indicate that our method maintains consistent performance across these varying hyperparameter configurations.
@@
Figure 4: Analysis of the performance growth between $S^{2}$ and FedAvg under varying scaling factors of NLIR and FGMA
Varying Client Scales. We assessed performance with different numbers of client partitions: 5, 10 and 20 clients. We compared $S^{2}$ with other methods, including FedAvg , FedProx , and FGSSL. The outcomes, presented in Fig. 5, demonstrate that $S^{2}$ consistently delivers reliable results regardless of the number of partitions.
@@
Figure 5: Analysis of performance under different client numbers.
Overall, $S^{2}$ exhibits strong stability and adaptability across varying hyperparameter settings and client partition configurations on both the Cora and Citeseer datasets. These findings confirm that $S^{2}$ not only sustains its effectiveness under diverse conditions but also adapts seamlessly to varying client scales, demonstrating its suitability for real-world subgraph-FL scenarios.
Q4: Does the effectiveness of NLIR help alleviate the impact of SIS decline in subgraph-FL?
Through experiments, we observe that as the number of clients increases, the SIS consistently decreases, indicating greater structural divergence across local graphs. Notably, the performance gains brought by NLIR become more pronounced under these conditions. This trend, illustrated in Fig. 6, highlights the targeted effectiveness of NLIR in mitigating the adverse effects of reduced SIS in subgraph-FL settings.
Figure 6: Performance improvement using the NLIR method in relation to the variation of the SIS across different client scales.
@@ 6 Conclusion
In this paper, we identify two critical challenges in subgraph-FL from the perspective of graph signal propagation: label signal disruption and spectral client drifts. To overcome these obstacles, we propose two key strategies: NLIR and FGMA. NLIR constructs a global repository of structurally representative nodes with rich label influence to deal with the LSD issue, while FGMA promotes a generic signal propagation paradigm, mitigating spectral client drift. By integrating these strategies, $S^{2}$ effectively addresses both the spatial and spectral challenges in FGL. Extensive experiments on various datasets validate that $S^{2}$ significantly boosts global generalizability.

### v3

@@ 5.2 Experiment Results
In this section, we comprehensively evaluate the proposed $S^{2}$FGL by addressing the following questions:
  • Q1: How does $S^{2}$FGL perform compared to existing FL and FGL methods in subgraph-FL?
  • Q2: What is the impact of each component of $S^{2}$FGL on the overall performance?
  • Q3: Does $S^{2}$FGL maintain consistent performance across different hyperparameter and client scales?
  • Q4: Do NLIR and FGMA mitigate the effects of label signal disruption and spectral heterogeneity?
Q1: How does $S^{2}$FGL perform compared to existing FL and FGL methods in subgraph-FL?
We present the results of node classification tasks across various FGL scenarios using multiple graph datasets, and we summarize the final average test accuracy in Tab. 1. It demonstrates that our proposed $S^{2}$FGL consistently outperforms all baseline approaches across all six datasets. This superiority highlights the effectiveness of $S^{2}$FGL.
Q2: What is the impact of each component of $S^{2}$FGL on the overall performance?
To evaluate the individual contributions of NLIR and FGMA strategies within the $S^{2}$FGL framework, we conducted ablation experiments on the Cora and Citeseer datasets. In this study, we removed each component to assess its impact on the overall performance. The results are presented in Tab. 2, which demonstrate that both NLIR and FGMA independently contribute to the overall performance.
@@
Table 2: Ablation study on key components of $S^{2}$FGL.
Q3: Does $S^{2}$FGL maintain consistent performance across different hyperparameter and client scales?
We evaluated the stability and adaptability of our proposed framework $S^{2}$FGL under varying hyperparameter configurations of and different client scales.
Varying Hyperparameters of NLIR and FGMA. For NLIR, we test hyperparameter sets of 100, 50, 10, and 1. For FGMA, the settings are 0.01, 0.05, 0.5, and 1. The results in Fig. 4 indicate that our method maintains consistent performance across varying hyperparameter configurations.
@@
Figure 4: Analysis of the performance growth between $S^{2}$FGL and FedAvg under different hyperparameters of NLIR and FGMA
Varying Client Scales. We assessed performance with different client scales: 5, 10 and 20. Specifically, we compared $S^{2}$FGL with other FL and FGL baselines, including FedAvg, FedProx, and FGSSL. The results in Fig. 5 demonstrate that $S^{2}$FGL consistently delivers reliable results regardless of the client scales.
@@
Figure 5: Analysis of performance under different client scales.
Overall, $S^{2}$FGL exhibits strong stability and adaptability across varying hyperparameter settings and client partition configurations on both the Cora and Citeseer datasets. These findings confirm that $S^{2}$FGL not only sustains its effectiveness under diverse conditions but also adapts seamlessly to varying client scales, demonstrating its suitability for real-world subgraph-FL scenarios.
Q4: Do NLIR and FGMA mitigate the effects of label signal disruption and spectral heterogeneity?
In Fig. 1, our experiments demonstrate that the SIS score in subgraph scenarios declines compared to centralized training, while spectral heterogeneity exists among clients. Here, we further verify the targeted effectiveness of NILR and FGMA with respect to these two issues, respectively. First, in Fig. 6 (a), we investigate the relationship between the performance gain brought by NILR to FedAvg and the change in the SIS score. The results show that this method achieves higher performance when semantic signals are more limited, thereby confirming its effectiveness and targeted nature. Second, in Fig. 6 (b), we examine the relationship between the performance improvement of FGMA for FedAvg and spectral heterogeneity, measured by the average KL divergence between the eigenvalue distributions of different clients. Experimental results show that greater spectral heterogeneity corresponds to larger performance gains, confirming the effectiveness of our method. Specifically, green indicates the performance gain of the proposed method relative to FedAvg, blue denotes variations in SIS, and orange corresponds to spectral heterogeneity.
(a) NILR
(b) FGMA
Figure 6: Analysis of the targeted effectiveness of NILR and FGMA. (a) The performance gain from NILR increases as the semantic signal becomes limited. (b) The performance improvement from FGMA grows with higher spectral heterogeneity.
@@ 6 Conclusion
In this paper, we identify and empirically demonstrate two phenomena in subgraph-FL from both spatial and spectral perspectives of graph signal propagation: label signal disruption and spectral heterogeneity. These phenomena pose challenges of poor semantic knowledge and spectral client drift. To address these challenges, we propose two key strategies: NLIR and FGMA. NLIR selects structurally and semantically representative nodes and constructs a global repository accordingly. By injecting semantic information from the repository into local training, it alleviates the poor semantic knowledge caused by label signal disruption. In addition, FGMA aligns and feature projections in both the high- and low-frequency reconstructed graph spectra, thereby promoting a generic signal propagation paradigm and mitigating client drifts under spectral heterogeneity. By integrating these strategies, $S^{2}$FGL effectively tackles both spatial and spectral challenges in subgraph-FL. Extensive experiments on multiple datasets demonstrate that $S^{2}$FGL significantly enhances global generalizability.